A bank loan money at 7% interest for five years to stew the long will be repaid in one lump sum payment of 3360 612 how much does tomorrow
Answer #1

FV = PV (1+r)^n

Where r is int rate per anum

n is time period

FV = PV (1+r)^n

3360612 = PV (1+0.07)^5

3360612 = PV (1.07)^5

3360612 = PV * 1.4026

PV = 3360612 / 1.4026

= $ 2,396,070

Pls comment, if any further assistance is required.